# MoDGS: Dynamic Gaussian Splatting from Causually-captured Monocular Videos

### v1

@@ 4.3. Comparison with Baselines
The qualitative results on the DyNeRF and Nvidia datasets are shown in Fig. 5. Other qualitative results on our MCV dataset are shown in Fig. 6. The quantitative results on the Nvidia and DyNeRF datasets are shown in Table 1 and Table 2 respectively.
@@
Synthesizing novel views from a casually captured monocular video is a challenging task. As shown in Fig. 5, though baseline methods achieve impressive results on these benchmarks with ”teleporting camera motions”, these methods fail to correctly reconstruct the 3D geometry of the dynamic scenes and produce obvious artifacts on both dynamic foreground and static background. The main reason is that the monocular camera is almost static and does not provide enough multiview consistency to reconstruct high-quality 3D geometry for novel view synthesis. In the third row of Fig. 5, SC-GS (Huang et al., 2024) fails to reconstruct the dynamic foreground subject because SC-GS has an initialization process that treats the whole scene as a static scene and trains on the scene for a number of steps. When the foreground subject is moving with a large motion (like skating from left to right), it would be ignored by the static scene initialization and then we fail to reconstruct in the subsequent steps.
@@
In comparison, our method relies on a 3D-aware initialization which provides a strong basis for the subsequent optimization. Meanwhile, our ordinal depth loss enables the 3D prior from the single-view depth estimator for an accurate reconstruction of the dynamic scenes. The quantitative results on both Table 2 and Table 1 also show that our method achieves the best performances in all metrics on both datasets. Note that there are still some artifacts on occlusion boundaries because the input monocular camera is almost static and these regions are not visible in our input videos.
Figure 7. Visualization of rendering results using and without using our 3D-aware initialization.
Figure 8. Visualization of the rendered depth and RGB images using our ordinal depth loss and the Pearson correlation loss.
4.4. Ablation studies
@@
We conduct ablation studies with our initialization and depth loss on the DyNeRF (Li et al., 2022) dataset to demonstrate their effectiveness. The qualitative results are shown in Fig. 7 and Fig. 8 while the quantitative results are shown in Table 3. We also provide additional results in the supplementary material to demonstrate MoDGS achieves much better rendering quality than simple depth warping.
@@
Table 3. Ablation studies with the 3D-aware initialization (”3D-aware Init”) and Depth Loss on the DyNeRF (Li et al., 2022) dataset. ”Ordinal” means the ordinal depth loss while ”Pearson” means the Pearson correlation loss.
@@
4.4.1. 3D-aware initialization
4.4.2. Ordinal depth loss
To demonstrate the effectiveness of our proposed ordinal depth, we experiment with two baseline settings, removing the depth loss and utilizing the Pearson correlation as the depth loss. As shown in Fig. 8, when there is no depth loss, the reconstructed 3D geometry contains much noise and obvious artifacts exist in the scene. Adopting the Pearson depth loss improves the quality by linear correlating the rendered depth maps and input depth maps but still produces noisy depth inside a region. In comparison, our ordinal depth loss enables a smooth reconstruction of the depth map in the interior while maintaining sharp edges at boundaries. Thus, the proposed ordinal depth loss enables a more robust reconstruction of the dynamic scene.
4.5. Limitations
Though our method can conduct dynamic view synthesis from casually captured monocular videos, the task is still extremely challenging. One limitation is that our method can only reconstruct the visible 3D parts but cannot imagine the unseen parts, which leads to artifacts when rendering novel view videos on these unseen parts. Incorporating recent 3D-related diffusion generative models (Chung et al., 2023; Liu et al., 2023b; Long et al., 2023) could be a promising direction to solve this problem, which we leave for future works. Another limitation is that the current training time is comparable with existing DVS methods, which could take several hours for a single scene. How to efficiently reconstruct the dynamic field would be an interesting and promising future research topic. Meanwhile, when the camera is completely static, our method strongly relies on the single-view depth estimator to estimate the 3D depth maps. Though existing single-view depth estimators (Fu et al., 2024; Ke et al., 2023; Yang et al., 2024; Bhat et al., 2023) are trained on large-scale datasets and predict reasonable depth maps for most cases, these depth estimators may fail to capture some details which degenerate the quality.
5. Conclusion
@@
Appendix A Supplementary Materials
A.1. Real Depth Recovery
The depth prediction of GeoWizard (Fu et al., 2024) is normalized depth values in [0,1] and we have to transform them into real depth values. We follow their official implementation to estimate a scaling factor and an offset value on the normalized normal maps by minimizing the normal maps estimated from the transformed real depth values and the estimated normal maps from GeoWizard. The optimization process takes just several seconds. Note that even after this normalization, the depth maps of different timesteps still differ from each other in scale.
Figure 9. We provide a pair of images on the iPhone (Gao et al., 2022) dataset. We draw two correspondences in the same color and their corresponding epipolar lines that are computed from the provided camera poses. The epipolar lines deviate far from the correspondences, which demonstrate that the camera poses are not accurate enough.
A.2. Pose Errors in the iPhone dataset
@@
A.3. Difference from Depth Map Warping
Figure 10. Comparison of our renderings and depth warping. Our method accumulates information among different timesteps and thus is able to render more completed images.
MoDGS learns a set of 3D Gaussians in a canonical space and a deformation field to transform it to an arbitrary timestep. This means that MoDGS is able to accumulate information among different timesteps to reconstruct a more completed scene than just using a single-view depth estimation. We show the difference between our renderings and just warping the training view using the estimated single-view depth map in Fig. 10. As we can see, MoDGS produces more completed reconstruction on contents that are not visible on this timestep. Meanwhile, MoDGS rectifies the single-view depth maps to be more accurate so the rendering quality is much better.
A.4. Pearson Depth Loss and Scale-shift Invariant Depth Loss
Previous works (Li et al., 2021; Liu et al., 2023a) use scale and shift invariant depth loss by minimizing the $L2$ distance of normalized ground truth depth map $\text{Norm}(D_{t})$ and normalized rendered depth map $\text{Norm}(\hat{D_{t}})$
(6) $\ell_{\text{depth}}=\|\text{Norm}(D_{t})-\text{Norm}(\hat{D}_{t})\|$
@@ A.4. Pearson Depth Loss and Scale-shift Invariant Depth Loss
(7) $\left(\cfrac{D_{t}-u_{D_{t}}}{\sigma_{D_{t}}}-\cfrac{D_{t}-u_{\hat{D_{t}}}}{\sigma_{\hat{D_{t}}}}\right)^{2}$
@@
where $D_{t}$ is the predicted depth prior, $\sigma_{D_{t}}$ and $u_{D_{t}}$ denotes the standard deviation and means respectively. The second term and third term are constant. The first term is a simple transformation of the Pearson correlation coefficient. Thus, minimizing the $L2$ distance is equivalent to maximizing the coefficient.

### v2

@@ 4.2 Comparison with Baselines
The qualitative results on the DyNeRF and Nvidia datasets are shown in Fig. 5. Other qualitative results on our MCV dataset are shown in Fig. 6. The quantitative results on the Nvidia and DyNeRF datasets are shown in Table 1. Supplementary videos contain more comparison results.
@@
Synthesizing novel views from a casually captured monocular video is a challenging task. As shown in Fig. 5, though baseline methods achieve impressive results on these benchmarks with “teleporting camera motions”, these methods fail to correctly reconstruct the 3D geometry of the dynamic scenes and produce obvious artifacts on both dynamic foreground and static background. The main reason is that the monocular camera is almost static and does not provide enough multiview consistency to reconstruct high-quality 3D geometry for novel view synthesis. In the third row of Fig. 5, SC-GS (Huang et al., 2024) fails to reconstruct the dynamic foreground subject because SC-GS has an initialization process that treats the whole scene as a static scene and trains on the scene for a number of steps. When the foreground subject is moving with a large motion (like skating from left to right), it would be ignored by the static scene initialization and then we fail to reconstruct in the subsequent steps.
@@
In comparison, our method relies on a 3D-aware initialization which provides a strong basis for the subsequent optimization. Meanwhile, our ordinal depth loss enables the 3D prior from the single-view depth estimator for an accurate reconstruction of the dynamic scenes. The quantitative results in Table 1 also show that our method achieves the best performances in all metrics on both datasets. Note that there are still some artifacts on occlusion boundaries because the input monocular camera is almost static and these regions are not visible in our input videos.
4.3 Ablation studies
@@
We conduct ablation studies with our initialization and depth loss on the DyNeRF (Li et al., 2022) dataset to demonstrate their effectiveness. The qualitative results are shown in Fig. 7 and Fig. 8 while the quantitative results are shown in Table 2. In the appendix, we also provide a comparison with depth warping, robustness to depth noises, a discussion on video depth, a comparison with other depth ranking losses, and a discussion about alpha values.
Figure 7: Visualization of rendering results using and without using our 3D-aware initialization.
4.3.1 3D-aware initialization
4.3.2 Ordinal depth loss
To demonstrate the effectiveness of our proposed ordinal depth, we experiment with two baseline settings, removing the depth loss and utilizing the Pearson correlation as the depth loss. As shown in Fig. 8, when there is no depth loss, the reconstructed 3D geometry contains much noise and obvious artifacts exist in the scene. Since we perform the 3D-aware initialization on this experiment to ensure a fair comparison, the results without depth loss still show reasonable rendered depth maps. Adopting the Pearson depth loss improves the quality by linear correlating the rendered depth maps and input depth maps but still produces noisy depth inside a region. In comparison, our ordinal depth loss enables a smooth reconstruction of the depth map in the interior while maintaining sharp edges at boundaries. Thus, the proposed ordinal depth loss enables a more robust reconstruction of the dynamic scene.
Figure 8: Visualization of the rendered depth and RGB images using our ordinal depth loss and the Pearson correlation loss.
@@ 4.3.2 Ordinal depth loss
Table 2: Ablation studies with the 3D-aware initialization (“3D-aware Init”) and Depth Loss on the DyNeRF (Li et al., 2022) dataset. “Ordinal” means the ordinal depth loss while “Pearson” means the Pearson correlation loss.
@@
4.4 Limitations
Though our method can conduct dynamic view synthesis from casually captured monocular videos, the task is still extremely challenging. One limitation is that our method can only reconstruct the visible 3D parts but cannot imagine the unseen parts, which leads to artifacts when rendering novel view videos on these unseen parts. Incorporating recent 3D-related diffusion generative models (Chung et al., 2023a; Liu et al., 2023b; Long et al., 2023) could be a promising direction to solve this problem, which we leave for future works. Another limitation is that the current training time is comparable with existing DVS methods, which could take several hours for a single scene. How to efficiently reconstruct the dynamic field would be an interesting and promising future research topic. Meanwhile, when the camera is completely static, our method strongly relies on the single-view depth estimator to estimate the 3D depth maps. Though existing single-view depth estimators (Fu et al., 2024; Ke et al., 2023; Yang et al., 2024; Bhat et al., 2023) are trained on large-scale datasets and predict reasonable depth maps for most cases, these depth estimators may fail to capture some details which degenerate the quality. Meanwhile, specular reflection can not be robustly and well handled, which is also an extremely challenging problem in the monocular setting, which we leave for future works.
5 Conclusion
@@
Appendix A Appendix
A.1 Implementation Details
We implement our MoDGS with PyTorch. To initialize the deformation field, we train it with 20k steps as stated in Sec. 3.2. Subsequently, we jointly train the 3D Gaussians and the deformation field with the rendering loss and the ordinal depth loss for another 20k steps. In Sec. 3.2, the flow is computed in evenly sampled key frames($1/5$). And the downsampling voxel size for Gaussian initialization is $0.004^{3}$ (scenes are normalized to $[-1,1]^{3}$ ). For the outer optimization loop and rendering loss, we exactly follow the original 3DGS. And we use Gaussian centers to render depth  (Yang et al., 2023b). We adopt an Adam optimizer for optimization. The learning rate for 3D Gaussians exactly follows the official implementation of 3D GS (Kerbl et al., 2023), while the learning rate of the deformation network undergoes exponential decay from 1e-3 to 1e-4 in initialization and from 1e-4 to 1e-6 in the subsequent optimization. We set $\alpha=100$ for $\ell_{\text{ordinal}}$. The weight of our depth order loss is 0.1. When computing depth ordinal loss, we first normalize the depth range to [0,1] and we only consider the depth pair with a difference larger than 0.02 for loss computation. The whole training takes around 3.5 hours to converge (2 hours for the initialization and 1.5 hours for the subsequent optimization) on an NVIDIA RTX A6000 GPU, which uses about 14G memory. The rendering speed of MoDGS is about 75 FPS.
A.2 Real Depth Recovery
The depth prediction of GeoWizard (Fu et al., 2024) is normalized depth values in [0,1] and we have to transform them into real depth values. We follow their official implementation to estimate a scaling factor and an offset value on the normalized normal maps by minimizing the normal maps estimated from the transformed real depth values and the estimated normal maps from GeoWizard. The optimization process takes just several seconds. Note that even after this normalization, the depth maps of different timestamps still differ from each other in scale.
Figure 9: We provide a pair of images on the iPhone (Gao et al., 2022) dataset. We draw two correspondences in the same color and their corresponding epipolar lines that are computed from the provided camera poses. The epipolar lines deviate far from the correspondences, which demonstrate that the camera poses are not accurate enough.
A.3 Pose Errors in the iPhone dataset
@@
A.4 Difference from Depth Map Warping
Figure 10: Comparison of our renderings and depth warping. Our method accumulates information among different timestamps and thus is able to render more completed images.
MoDGS learns a set of 3D Gaussians in a canonical space and a deformation field to transform it to an arbitrary timestamp. This means that MoDGS is able to accumulate information among different timestamps to reconstruct a more completed scene than just using a single-view depth estimation. We show the difference between our renderings and just warping the training view using the estimated single-view depth map in Fig. 10. As we can see, MoDGS produces more completed reconstruction on contents that are not visible on this timestamp. Meanwhile, MoDGS rectifies the single-view depth maps to be more accurate so the rendering quality is much better.
A.5 Pearson Depth Loss and Scale-shift Invariant Depth Loss
In this section, we will prove that the Pearson Depth loss is equivalent to Scale-Shift Invariant loss. Previous works (Li et al., 2021; Liu et al., 2023a) use scale and shift invariant depth loss by minimizing the $L2$ distance of normalized ground truth depth map $\text{Norm}(D_{t})$ and the normalized rendered depth map $\text{Norm}(\hat{D_{t}})$
$\ell_{\text{depth}}=\|\text{Norm}(D_{t})-\text{Norm}(\hat{D}_{t})\|$ (5)
@@ A.5 Pearson Depth Loss and Scale-shift Invariant Depth Loss
$\left(\cfrac{D_{t}-u_{D_{t}}}{\sigma_{D_{t}}}-\cfrac{D_{t}-u_{\hat{D_{t}}}}{\sigma_{\hat{D_{t}}}}\right)^{2}$ (6)
@@
where $D_{t}$ is the predicted depth prior, $\sigma_{D_{t}}$ and $u_{D_{t}}$ denotes the standard deviation and means respectively. Since it is based on the input, The second term is a constant. For the third term, we substitute the $\sigma_{\hat{D_{t}}}$ using its definition: $\sigma_{\hat{D_{t}}}=\sqrt{\frac{\sum_{i=1}^{N}\left(\hat{D_{t}^{i}}-u_{\hat{D_{t}}}\right)^{2}}{N}}$. The sum of the third term over all pixels in a depth map is: $\Sigma_{i}^{N}(\left(\cfrac{\hat{D_{t}^{i}}-u_{\hat{D_{t}}}}{\sigma_{\hat{D_{t}}}}\right)^{2})=N$, Where $N$ is the total number of pixels.
@@
We can see that both the second term and third term are constant. The first term is a simple transformation of the Pearson correlation coefficient. Thus, minimizing the $L2$ distance is equivalent to maximizing the coefficient.
A.6 Results on the Davis dataset
A.6.1 Preprocessing Procedure
On the Davis dataset, we adopt the preprocessing procedure outlined in the concurrent method, Shape-of-Motion (Wang et al., 2024b), to obtain these camera poses. Specifically, we first run UniDepth (Piccinelli et al., 2024) to acquire camera intrinsics and depth maps. Then, we utilize a SLAM method, Driod-SLAM (Teed & Deng, 2021), with the depth maps from UniDepth as input to solve for the camera poses.
A.6.2 Comparison with Shape-of-Motion
Figure 11: We show a visual comparison of rendered novel view images and depth maps between Shape-of-Motion(Wang et al., 2024b) and MoDGS. The rendering quality is comparable but our method shows more details in some cases.
We provide an experiment to compare Shape-of-Motion with our method. In our experiments, we train shape-of-motion on the Davis dataset following their default training setting and use the same preprocessing strategy like depth estimation to train MoDGS. The visual comparison is shown in Fig. 11, which demonstrates that our method shows comparable results as Shape-of-Motion but renders more details in some cases. We also present the quantitative comparison results on the Nvidia dataset  (Yoon et al., 2020) in Table 3.
Table 3: Comparison with Shape-of-Motion(SoM) on Nvidia dataset
Balloon2-2 Jumping Truck-2 Skatting2
Method PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
SoM 19.52 0.5096 0.2790 20.32 0.6771 0.2560 22.53 0.7414 0.1853 23.51 0.7861 0.1658
Ours 20.47 0.5275 0.2408 22.18 0.7075 0.2125 23.69 0.7455 0.1551 25.64 0.7996 0.1518
A.7 Discussion about alpha value in ordinal depth loss
After roughly aligning the depth maps, we normalize the scene to $[-1,1]^{3}$ and empirically set alpha to 100. We present further ablations in Table 4, where decreasing alpha leads to worse results while increasing alpha yields comparable results. The reason is that large alpha values compel the loss to mainly concentrate on the ordinality rather than the differences in depth values.
Table 4: Quantitative results for different alpha values.
Setting PSNR SSIM LPIPS
alpha=10 21.05 0.7575 0.2370
alpha=40 21.39 0.7813 0.1916
alpha=100(Ours) 23.23 0.8083 0.1592
alpha=1000 23.16 0.8058 0.1715
A.8 Discussion about the Robustness to the Depth noises
To show the robustness of our Ordinal Depth Loss to depth noise, we manually add Gaussian noises to depth maps and use the noisy depth maps as supervision signals. The metric results of them are shown in Table  5. (on the “flame_steak” scene), which shows our method is robust to small depth noises.
Table 5: Quantitative results with different levels of added Gaussian noise $Norm(\mu,\sigma)$
Setting PSNR SSIM LPIPS
$Norm(0,0.20)$ 23.18 0.8090 0.1667
$Norm(0,0.05)$ 23.35 0.8100 0.1633
Ours 23.23 0.8083 0.1592
A.9 Per-Scene breakdown results on the DyNeRF  (Li et al., 2022) dataset and the Nvidia (Yoon et al., 2020) dataset
In Tables 6 and Table 7, we provide a breakdown of the metrics for different scenes in the DyNeRF and Nvidia datasets, respectively.
Table 6: Per-scene results on the DyNeRF dataset.
Method cut_roasted_beef sear_steak coffee_martini
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
HexPlane 16.76 0.5382 0.5054 16.89 0.5897 0.5049 13.26 0.4049 0.5835
SC-GS 20.69 0.7414 0.2625 21.23 0.7870 0.2188 19.02 0.7124 0.2151
D-GS 22.20 0.7808 0.1931 23.56 0.8101 0.1773 19.23 0.7013 0.2270
Ours 23.98 0.8221 0.1438 23.53 0.8126 0.1642 21.37 0.7962 0.1473
Method cook_spinach flame_steak flame_salmon_1
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
HexPlane 16.95 0.7286 0.2223 16.97 0.7528 0.2543 11.16 0.3417 0.6382
SC-GS 16.70 0.7377 0.2117 17.31 0.7532 0.2527 17.65 0.6834 0.2253
D-GS 17.20 0.7195 0.2329 16.62 0.7523 0.2559 18.48 0.7038 0.2166
Ours 22.40 0.7823 0.1728 23.23 0.8083 0.1592 21.33 0.8038 0.1399
Table 7: Per-scene results on the Nvidia dataset.
Method Playground DynamicFace-2 Balloon1-2 Umbrella
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
HexPlane 13.30 0.1641 0.5240 9.92 0.1873 0.6091 15.48 0.2793 0.5342 18.70 0.2238 0.4427
SC-GS 12.60 0.2059 0.5063 9.75 0.3312 0.4307 16.63 0.4412 0.3656 18.71 0.3501 0.3663
D-GS 12.62 0.2009 0.5116 10.36 0.3277 0.4563 10.36 0.3277 0.4563 19.16 0.3381 0.3300
Ours 13.38 0.2343 0.4542 12.40 0.4118 0.2442 16.93 0.4371 0.3059 19.47 0.3246 0.3004
Method Balloon2-2 Skatting2 Truck Jumping
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
HexPlane 18.42 0.3159 0.4881 21.39 0.6403 0.3983 20.85 0.5687 0.3653 19.28 0.5605 0.4431
SC-GS 18.28 0.3893 0.3671 23.07 0.7186 0.2002 21.48 0.6374 0.2078 20.17 0.6698 0.2343
D-GS 18.57 0.3941 0.3768 24.47 0.7582 0.2073 21.38 0.6321 0.2142 21.27 0.6604 0.2700
Ours 20.47 0.5275 0.2408 25.64 0.7996 0.1518 23.69 0.7455 0.1551 22.18 0.7075 0.2125
A.10 Comparison with depth ranking Loss
We provide a quantitative comparison between our ordinal depth loss and other depth ranking losses (Chen et al., 2016; Pavlakos et al., 2018) in Table 8. In this experiment, we replace our ordinal depth loss with their depth ranking loss while keeping all other configurations fixed. The results demonstrate that our ordinal depth loss significantly outperforms their approach across all scenes. This highlights the effectiveness of our method in capturing depth information more accurately than the existing depth ranking losses.
In Fig. 12, we present a visual comparison of rendered depth maps using depth ranking loss and ordinal depth loss. Our depth values are more smooth and well-distributed while depth values learned from rank depth loss show high contrast and more noise. The reason is that depth ranking loss encourages large contrast to satisfy the given depth order and is not robust to the noises in the input depth.
Figure 12: A visual comparison of rendered depth maps between using depth ranking loss (Wang et al., 2024b) and our ordinal depth loss.
Table 8: Comparison with depth ranking loss  (Chen et al., 2016; Pavlakos et al., 2018),denoted as $DRL$ for simplification.
sear_steak cut_beef coffee_martini
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
$DRL$ 18.22 0.7390 0.2228 17.07 0.7162 0.2283 14.09 0.7465 0.1902
Ours 23.53 0.8126 0.1642 23.98 0.8221 0.1438 21.37 0.7962 0.1473
A.11 ablation using depth from Video-depth estimation methods
To show that our method is also compatible with recent video depth methods, we conduct an ablation experiment using the depth maps predicted by the video depth estimator  (Shao et al., 2024) and different types of depth loss as supervision. As shown in Table 9, the video depth estimator improves the results, and our ordinal depth loss still consistently improves the results.
Table 9: Video depth ablation. $VD+P$ refers to supervised using video depth maps with pearson depth loss, $VD+O$ refers to supervised using video depth maps with ordinal depth loss.
sear_steak cut_beef coffee_martini
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
$VD+P$ 23.43 0.7983 0.1732 23.19 0.8134 0.1472 20.01 0.7664 0.1647
$VD+O$ 23.48 0.8091 0.1711 23.61 0.8103 0.1532 21.55 0.8001 0.1488
A.12 Results of NeRF-based methods using depth supervision
We present the results of HexPlane (Cao & Johnson, 2023) when supervised with depth loss in Table 10. Using depth maps for supervision only leads to a slight improvement for HexPlane and MoDGS still outperforms HexPlane by a large margin.
Table 10: Metric results when HexPlane  (Cao & Johnson, 2023) is incorporated with depth supervision, denoted as “HexPlane+D”.
flame_steak cut_beef coffee_martini
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
HexPlane 16.97 0.7528 0.2543 16.76 0.5382 0.5054 13.26 0.4049 0.5835
HexPlane+D 17.21 0.6156 0.4646 18.82 0.6344 0.414 15.78 0.4708 0.5106
Ours 23.23 0.8083 0.1592 23.98 0.8221 0.1438 21.37 0.7962 0.1473
A.13 Different random initialization methods
In the ablation study of our paper, we use the Kaiming initialization as the random initialization. We additionally provide ablations using zero deformation initialization of the deformation field in Table 11(flame_steak scene of DyNeRF). In this experiment, we still initialize the canonical Gaussian spaces but adopt a different initialization strategy for the deformation field MLP.
Table 11: Quantitative comparison results for different initialization methods
Method PSNR SSIM Lpips
ZeroDeformationInit 22.34 0.7906 0.2151
KaimingInit 22.24 0.7744 0.2036
3D-AwareInit (Ours) 23.23 0.8083 0.1592
A.14 Length of videos that MoDGS can handle
Currently, we can handle videos up to 10 seconds. Scaling to longer videos($\sim$30 seconds) is possible if we use a larger deformation MLP network and train for more iterations.